According to a Pew Research Center study, in May 2011, 31% of all American adults had a smart phone (one which the user can use to read email and surf the Internet). A communications professor at a university believes this percentage is higher among community college students. She selects 352 community college students at random and finds that 129 of them have a smart phone. In testing the hypotheses:

H0: p = 0.31 versus

Ha: p > 0.31,

she calculates the test statistic as z = 2.2911.

Then the p‑value = .

(Please round your answer to four decimal places.)
Answer #1

solution:

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =0.31

Ha : p > 0.31

\hat p = x / n = 129/352=0.37

P0 = 0.31

1 - P0 = 1-0.31=0.69

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 2.2911 given

p value =p(z>2.29)=1-p(z<2.29)=1-0.9890=0.0110

p value =0.0110